# Learning Exposure Correction in Dynamic Scenes

### v2

@@ 3.3 Dual-stream Illumination Construction
To learn adaptive representations for both underexposure and overexposure, we exploit dual-stream illumination construction in our model. From the Retinex theory, the reference frame $\mathcal{I}_{t}$ can be typically decomposed into an illumination map $L$ and a reflectance map $\mathcal{R}$. However, this strategy can only be learned from underexposure conditions as $L$ falls into the values within [0,1]. Existing Retinex-based methods [4, 7] would increase the exposure of the overexposed inputs.
To suppress overexposure while enhancing underexposure, we propose a dual-stream mechanism to achieve illumination construction, sharing parameters with each other. Specifically, as shown in Fig. 6, we treat the antithetic exposure $\mathcal{I}^{inv}_{t}$ as the reverse frame of $\mathcal{I}_{t}$. We first encode $\{\mathcal{I}_{t},\mathcal{I}^{inv}_{t}\}$ to $z_{t},z^{inv}_{t}$ in the latent space. Then we use a U-Net [21] to learn the mapping from $\{z_{t},z^{inv}_{t}\}$ to $\{\tilde{z}_{t},\tilde{z}^{inv}_{t}\}$. Finally we decode $\{\tilde{z}_{t},\tilde{z}^{inv}_{t}\}$ to $\{L_{t},L^{inv}_{t}\}$ in the image space. By calculating with the learned and extended illumination map $L^{inv}$, we can obtain the enhanced image of overexposure $\mathcal{I}^{o}_{t}$. This process can be formulated as:
$\begin{split}\mathcal{I}^{u}_{t}&=\mathcal{I}_{t}\cdot L_{t}^{-1},\\ \mathcal{I}^{inv}_{t}&=1-\mathcal{I}_{t},\\ \mathcal{I}^{o}_{t}&=1-\mathcal{I}^{inv}_{t}\cdot(L_{t}^{inv})^{-1},\end{split}$ (5)
where $\mathcal{I}^{u}_{t}$ represents the enhanced frame of underexposure, $\cdot$ denotes element-wise multiplication.
@@ 3.5 Objectives
We incorporate a range of loss terms to facilitate the training of our model. We employ the commonly utilized pixel-wise charbonnier loss term, denoted as $\mathcal{L}_{pix}$, to measure the accuracy of pixel reconstruction. Based on Retinex theory, we present reflectance consistency loss $\mathcal{L}_{R}$ between paired generated frames and ground truths. To guide the network to estimate and smooth the dual illumination maps, we apply the total variation [24] loss term $\mathcal{L}_{tv}$. They can be formulated as:
$\begin{split}\mathcal{L}_{R}&=||\mathcal{T}(\mathcal{O}_{t})-\mathcal{T}(\mathcal{I}_{t}^{gt})||^{2}_{F},\\ \mathcal{L}_{tv}&=\sum_{c}[{(\partial_{x/y}L_{t})}^{2}+{(\partial_{{x/y}}L_{t}^{inv})}^{2}],\end{split}$ (7)
where all channels (c) of all pixels are summed, $\partial_{x/y}$ are partial derivatives in horizontal and vertical directions in the image space. In addition, In order to ensure exposure continuity between video frames, we introduce amplitude consistency loss term $\mathcal{L}_{amp}$. It is formulated as:
@@
where $X_{(\cdot)}$ represents the output or ground truth normal exposure image in Fourier space. The overall function can be written as:
$\begin{split}\mathcal{L}_{total}&=\lambda_{1}\mathcal{L}_{pix}+\lambda_{2}\mathcal{L}_{R}+\lambda_{3}\mathcal{L}_{tv}+\lambda_{4}\mathcal{L}_{amp},\end{split}$ (9)
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are weight hyperparameters.
@@ 4 Experiments
Figure 8: Qualitative comparisons of the multi-frame results with state-of-the-art methods on the DIME dataset.
@@ 4.1 Implementation Details
We implement our framework with Pytorch on a single NVIDIA GeForce RTX 3090 GPU. We use $N=2$ consecutive frames without intervals as input. We divide the DIME dataset into training/validation/test sets according to the number of videos in 90:9:20. Then we use the training set to train our model with a total of 400,000 iters. The parameters of the network are optimized by the ADAM optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ The learning rate is 0.001 while the batch size is 8. The frames apply random horizontal and vertical flipping to augment the input data. We train on patches of size $256\times 256$ pixels. The weights for the terms in the loss function in Eq. (9) are $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}=1.0,1.0,0.01,100.0$. The additional experimental results are presented in the supplementary file.
@@ 4.2 Baselines
Since there is currently no related method for video exposure correction, we adopt several image exposure correction methods for comparison including MSEC [1], DRBN-ENC [10], ECLNet [12], FECNet [11], and LACT [2]. We also take low-level tasks adjacent to video exposure correction, including several methods in the fields of video low-light enhancement and video restoration, such as SMOID [13], SDSD [25], RVRT [17], and DIDNet [8].
@@
DIDNet [8] 16.20 0.8435 18.61 0.9022 17.41 0.8729
MSEC [1] 13.22 0.8071 18.01 0.8854 15.61 0.8462
DRBN-ENC [10] 15.69 0.8380 18.02 0.9203 16.83 0.8782
ECLNet [12] 13.16 0.7662 20.77 0.9105 16.88 0.8367
FECNet [11] 14.65 0.8461 22.52 0.9009 18.49 0.8729
LACT [2] 12.91 0.7688 17.62 0.8236 15.26 0.7962
VECNet (Ours) 18.63 0.8987 21.91 0.9379 20.27 0.9183
Table 1: Quantitative PSNR and SSIM results of different methods on Underexposed and Overexposed samples of the DIME dataset. A higher metric indicates better performance. The best and second results are marked in bold and underlined, respectively.
Method DIME LLIV-Phone
@@
SMOID [13] 5.43 35.32 6.76 6.06
SDSDNet [25] 4.35 18.04 4.07 8.99
RVRT [17] 4.17 20.81 5.75 7.07
DIDNet [8] 7.25 42.60 6.93 10.32
MSEC [1] 5.38 37.28 5.80 232.8
DRBN-ENC [10] 4.48 45.51 5.24 129.4
ECLNet [12] 4.40 23.31 5.38 141.0
FECNet [11] 4.52 29.11 5.26 105.5
LACT [2] 5.01 35.57 5.62 171.6
VECNet (Ours) 4.10 15.72 4.01 4.65
Table 2: Quantitative results of different methods in terms of NIQE and ALV. A lower metric indicates better performance. The best and second results are marked in bold and underlined, respectively.
@@ 4.3 Quantitative Evaluation
We demonstrate the superiority of our method and the impact of the DIME dataset through experiments in this section. We conduct quantitative experiments on the test set, which consists of 10 underexposed video pairs and 10 overexposed video pairs with a total of 3,824 frames, as described in Sec. 2. We adopt the following three standard metrics to evaluate the pixel-wise accuracy and the perceptual quality of our results: reference-based (i) Peak Signal-to-Noise Ratio (PSNR) and (ii) Structural Similarity Index Measure (SSIM) [30], non-reference based (iii) Natural Image Quality Evaluator (NIQE) [20] and (iv) Average Luminance Variance (ALV) [15] metrics. Table 1 reports the quantitative evaluation results of different methods. As exhibited in the table, our proposed VECNet outperforms all other methods in PSNR and SSIM metrics, demonstrating its superior performance in video exposure correction tasks.
Validation on unpaired videos. To evaluate the generalization ability of our method under challenging scenes with large-scale motions, we also conduct extensive experiments on the LLIV-Phone dataset [15]. Since there is no ground truth, all compared baselines are only tested on the dataset. Note that we report the results of our model trained on the DIME training set without further tuning or re-training on any of these datasets. Table 2 reports the quantitative results. Our method achieves the best results in terms of NIQE and ALV, demonstrating the stability when handling real-world videos with various motions.
Various kinds of device video enhancement. We collect a total of 30 abnormal exposure videos from three devices, 10 from each device. We test them with the compared methods pretrained on the DIME training set. As described in Table 5, our method achieves better results, displaying the effectiveness for more applications. We present the luminance curves of the enhanced video frames to verify the degree and continuity of the exposure, as shown in Fig. 11.
4.4 Qualitative Evaluation
We perform thorough qualitative evaluations on the DIME and LLIV-Phone datasets to assess the performance of our proposed method. We present single-frame and multi-frame results in Fig. 8 and Fig. 9, which indicates that VECNet delivers more natural and reasonable enhancement. In particular, low-light video enhancement methods mainly produce frames with unpleasing regions, leading to substantial loss of texture. Image-based exposure correction methods result in significant color deviation, which adversely affects the visual quality of the images. SMOID and RVRT produce frames with severe blocking artifacts. Moreover, image-based methods generate exposure inconsistent frames as the temporal information has not been well utilized to avoid flickering. After a comprehensive evaluation of the comparative results of different methods on two datasets, our proposed method exhibits excellent visual performance in terms of global brightness, color recovery, and detail while maintaining temporal stability without flickering artifacts and motion blur.
Model MPA DIC TSR PSNR SSIM
phase single dual stage-1 stage-2
(a) 18.20 0.7251
(b) 18.78 0.8728
(c) 19.12 0.9025
(d) 20.27 0.9183
Table 3: Ablation study for investigating the components of the specific modules. For the MPA module, we ablate the phase operation. For the DIC module, we ablate the single and dual streams.
@@ 4.4 Qualitative Evaluation
Figure 9: Qualitative comparisons of the multi-frame results with state-of-the-art methods on the LLIV-Phone dataset.
Method Other Same VECNet
SMOID [13] 0.20 0.15 0.65
SDSDNet [25] 0.25 0.20 0.55
RVRT [17] 0.05 0.15 0.80
DIDNet [8] 0.25 0.10 0.65
Table 4: User study score comparisons with baselines. “VECNet” is the percentage that our result is preferred, “Other” is the percentage that some other approach is preferred, “Same” is the percentage that the users have no preference.
@@
VECNet 3.44/19.01 4.49/25.83 3.49/8.37 3.81/17.74
Table 5: Quantitative results of different methods for multi-exposure videos captured from various devices and internet, evaluated by NIQE and ALV metrics.
@@ 4.5 User Study
We conduct a user study with 30 participants to evaluate the subjective perception of different methods. We select 20 testing videos from the DIME dataset. The videos are then enhanced using five video enhancement methods (SMOID, SDSDNet, RVRT, DIDNet, and VECNet) to perform pairwise comparisons. For a fair comparison, the user study is conducted in the same environment (room, display, and light). Participants are asked to choose: “baseline is better”, “VECNet is better”, or “no preference” in terms of brightness, naturalness, contrast, color, and temporal stability. In Table 4, our method is preferred by human subjects, suggesting that our results are more natural and realistic.
4.6 Ablation Study
We provide a series of ablation studies to evaluate the effectiveness of each component in the proposed method. Note that all experiments are conducted on the DIME dataset. The ablation results are shown in Table 3 and Fig. 10.
For the MPA module, we remove it and directly fuse the original unaligned multiple frames to learn reflection maps to generate the final result. It can be seen MPA brings expected quantitative performance improvement. The results of Fig. 10 (b) are more visually appealing with fewer noises and motion blurs than (a), which further demonstrates the effectiveness of MPA. For the DIC unit, we replace the symmetrical exposure stream with a single path setting. It can be seen the overexposure correction shows little improvement from Fig. 10 (b). For the TSR unit, we split the TSR unit into two separate stages. Comparing Fig. 10 (c) and (d), we can see that our two-stage strategy helps preserve more color and texture information.
Figure 10: Visualizations of ablation study on proposed module.
Figure 11: Lux curves of the video enhanced by different methods. Left is underexposure cases while right is overexposure cases.
@@ 5 Conclusion
We build the first high-quality paired video exposure correction dataset for dynamic real-world scenes with multi-exposures, camera and object motions, and precise spatial alignment. Based on the dataset, we propose a method by dealing with the underexposed and overexposed inputs in a dual-stream manner. By utilizing phase alignment and synthesis modules, the exposure of the videos is well corrected and restored. Experiments demonstrate that the proposed method outperforms several image/video methods in low-level tasks adjacent to video exposure correction.

### v3

@@ 3.3 Dual-stream Illumination Construction
To learn adaptive representations for both underexposure and overexposure situations, we exploit dual-stream illumination construction in our model. From the Retinex theory, the reference frame $\mathcal{I}_{t}$ can be typically decomposed into an illumination map $L$ and a reflectance map $\mathcal{R}$. However, this strategy can only be learned from underexposure conditions, as $L$ falls into the range of values within [0, 1]. Existing Retinex-based methods [4, 8] increase the exposure level of the overexposed inputs.
To suppress overexposure while enhancing underexposure, we propose a dual-stream mechanism to achieve illumination construction in both latent space and image level, rather than only in image level [26], sharing parameters with each other. Specifically, as shown in Fig. 6, we treat the antithetic exposure $\mathcal{I}^{inv}_{t}$ as the reverse frame of $\mathcal{I}_{t}$. We first encode $\{\mathcal{I}_{t},\mathcal{I}^{inv}_{t}\}$ to $z_{t},z^{inv}_{t}$ in the latent space. Then we use a U-Net [22] to learn the mapping from $\{z_{t},z^{inv}_{t}\}$ to $\{\tilde{z}_{t},\tilde{z}^{inv}_{t}\}$. Finally, we decode $\{\tilde{z}_{t},\tilde{z}^{inv}_{t}\}$ to $\{L_{t},L^{inv}_{t}\}$ in the image space. By calculating with the learned and extended illumination map $L^{inv}$, we can obtain the enhanced image of overexposure $\mathcal{I}^{o}_{t}$. This process can be formulated as:
$\begin{split}\mathcal{I}^{u}_{t}=\mathcal{I}_{t}\cdot L_{t}^{-1},\quad\mathcal{I}^{inv}_{t}=1-\mathcal{I}_{t},\quad\mathcal{I}^{o}_{t}=1-\mathcal{I}^{inv}_{t}\cdot(L_{t}^{inv})^{-1},\end{split}$ (6)
where $\mathcal{I}^{u}_{t}$ represents the enhanced frame of underexposure, and $\cdot$ denotes element-wise multiplication.
@@ 3.5 Objectives
We incorporate a range of loss terms to facilitate the training of our model. We employ the commonly utilized pixel-wise charbonnier loss term, denoted as $\mathcal{L}_{pix}$, to measure the accuracy of pixel reconstruction. To guide the network to estimate and smooth the dual illumination maps, we apply the total variation [27] loss term $\mathcal{L}_{tv}$. They can be formulated as:
$\begin{split}\mathcal{L}_{tv}&=\sum_{c}[{(\partial_{x/y}L_{t})}^{2}+{(\partial_{{x/y}}L_{t}^{inv})}^{2}],\end{split}$ (8)
where all channels (c) of all pixels are summed, $\partial_{x/y}$ are partial derivatives in horizontal and vertical directions in the image space. In addition, to ensure exposure continuity between video frames, we introduce the amplitude consistency loss term $\mathcal{L}_{amp}$, which can be formulated as:
@@
where $X_{(\cdot)}$ represents the frequency component of output $O_{t}$ or ground truth normal exposed image $I^{gt}_{t}$ in Fourier space. The overall function can be written as:
$\begin{split}\mathcal{L}_{total}&=\lambda_{1}\mathcal{L}_{pix}+\lambda_{2}\mathcal{L}_{tv}+\lambda_{3}\mathcal{L}_{amp},\end{split}$ (10)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are weight hyperparameters.
Figure 8: Qualitative comparisons of the multi-frame results with different methods on the DIME dataset.
@@ 4.1 Implementation Details
We implement our framework with Pytorch on a single NVIDIA GeForce RTX 3090 GPU. We use $N=2$ consecutive frames without intervals as input. We divide the DIME dataset into training, validation, testing sets according to the number of videos in 90:9:20. Then we use the training set to train our model for a total of 800,000 iterations. The parameters of the network are optimized by the ADAM optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. The learning rate is 0.001, while the batch size is 8. The frames apply random horizontal and vertical flipping to augment the input data. We train on patches of size $256\times 256$. The weights for the terms in the loss function in Eq. (9) are $\lambda_{1},\lambda_{2},\lambda_{3}=1.0,0.01,100.0$.
@@ 4.2 Baselines
Since there is currently no related method for video exposure correction, we adopt several image exposure correction methods for comparison, including MSEC [1], DRBN-ENC [11], ECLNet [13], FECNet [12], and LACT [2]. We also take low-level tasks adjacent to video exposure correction, including several methods in the fields of video low-light enhancement and video restoration, such as SMOID [14], SDSD [28], RVRT [18], and DIDNet [9].
Table 1: Quantitative PSNR and SSIM results of different methods on Underexposed and Overexposed samples of the DIME dataset. A higher metric indicates better performance. The best and second-best results are in bold and underlined.
@@
DIDNet [9] 17.89 0.8587 21.01 0.9022 19.45 0.8805
MSEC [1] 16.22 0.8271 18.01 0.8854 17.12 0.8563
ENC [11] 17.69 0.8580 20.12 0.9203 18.91 0.8892
ECLNet [13] 16.80 0.8062 19.56 0.8915 18.18 0.8489
FECNet [12] 17.61 0.8601 21.02 0.9109 19.32 0.8855
LACT [2] 18.01 0.8651 20.36 0.9153 19.19 0.8902
VECNet (Ours) 18.18 0.8687 22.04 0.9345 20.11 0.9016
Table 2: Quantitative results of different methods in terms of NIQE and ALV. A lower metric indicates better performance. The best and second-best results are in bold and underlined.
Method (Publication) DIME LLIV-Phone
@@
SMOID [14] (ICCV’19) 5.43 35.32 6.76 6.06
SDSDNet [28] (ICCV’21) 4.35 18.04 4.07 8.99
RVRT [18] (NeurIPS’22) 4.17 20.81 5.75 7.07
DIDNet [9] (ICCV’23) 7.25 42.60 6.93 10.32
MSEC [1] (CVPR’22) 5.38 37.28 5.80 232.8
DRBN-ENC [11] (CVPR’22) 4.48 45.51 5.24 129.4
ECLNet [13] (ACMMM’22) 4.40 23.31 5.38 141.0
FECNet [12] (ECCV’22) 4.52 29.11 5.26 105.5
LACT [2] (ICCV’23) 5.01 35.57 5.62 171.6
VECNet (Ours) 4.08 15.65 4.02 4.94
@@ 4.3 Quantitative Evaluation
We demonstrate the superiority of our method and the impact of the DIME dataset through experiments in this section. We conduct quantitative experiments on the test set, which consists of 10 underexposed video pairs and 10 overexposed video pairs with a total of 3,824 frames, as described in Sec. 2. We adopt the following three standard metrics to evaluate the pixel-wise accuracy and perceptual quality of our results: reference-based (i) Peak Signal-to-Noise Ratio (PSNR) and (ii) Structural Similarity Index Measure (SSIM) [33], non-reference-based (iii) Natural Image Quality Evaluator (NIQE) [21] and (iv) Average Luminance Variance (ALV) [16] metrics. Table 1 reports the quantitative evaluation results of different methods. As exhibited in the table, our proposed VECNet outperforms all other image-based and video enhancement and restoration-based methods in PSNR and SSIM metrics, demonstrating its superior performance in video exposure correction tasks.
@@
Figure 9: Qualitative comparisons of the multi-frame results with different methods on the LLIV-Phone dataset.
(a) VECNet (Ours)
(b) SMOID [14]
(c) SDSDNet [28]
(d) RVRT [18]
(e) DIDNet [9]
Figure 10: The results of five methods in the user study. In each histogram, the x-axis denotes the ranking index (1$\sim$5, 1 represents the highest), and the y-axis denotes the number of images in each ranking index.
Validation on unpaired videos. To evaluate the generalization ability of our method under challenging dynamic scenes, we also conduct extensive experiments on the LLIV-Phone dataset [16]. Since there is no ground truth, all compared baselines are only tested on the dataset. Note that we report the results of our model trained on the DIME training set without further tuning or re-training on any of these datasets. Table 2 reports the quantitative results. Our method achieves the best results in terms of NIQE and ALV, demonstrating stability when handling real-world videos with various motions.
Various kinds of device video enhancement. We collect a total of 30 abnormal exposure videos from three devices, 10 from each device. We test them with the compared methods pretrained on the DIME training set. As described in Table 3, our method achieves better results, displaying its effectiveness for more applications. In addition, we present the luminance curves of the enhanced video frames to verify the degree and continuity of the exposure, as shown in Fig. 12. VECNet behaves more stable, especially in some cases with drastic light changes (solid line).
Table 3: Quantitative results of different methods for multi-exposure videos captured from various devices and the internet, evaluated by NIQE and ALV metrics.
@@
VECNet 3.28/22.95 4.40/26.04 3.45/10.77 3.71/19.92
4.4 Qualitative Evaluation
We perform thorough qualitative evaluations on the DIME and LLIV-Phone datasets to assess the performance of our proposed method. We present single-frame and multi-frame results in Fig. 8 and Fig. 9, which indicates that VECNet delivers more natural and reasonable enhancement. In particular, low-light video enhancement methods mainly produce frames with unpleasing regions, leading to substantial loss of texture. Image-based exposure correction methods result in significant color deviation, which adversely affects the visual quality of the images. SMOID and RVRT produce frames with severe blocking artifacts. Moreover, image-based methods generate inconsistent exposure frames as the temporal information has not been well utilized to avoid flickering. After a comprehensive evaluation of the comparative results of different methods on two datasets, our proposed method exhibits excellent visual performance in terms of global brightness, color recovery, and detail while maintaining temporal stability without flickering artifacts and motion blur.
@@ 4.5 User Study
We conduct a user study with 20 participants to evaluate the subjective perceptions of different methods. We select 20 testing videos from the DIME dataset. The videos are then enhanced using 5 video enhancement methods (SMOID, SDSDNet, RVRT, DIDNet, and VECNet) to perform pairwise comparisons. For a fair comparison, we provide the original input videos and present 5 kinds of enhancement results, randomly changing the display order to prevent bias. The user study is conducted in the same environment (room, display, and light). Participants are asked to rate their video quality from best to worst. The human subjects are instructed to consider authenticity, exposure artifacts, texture contrast, realistic color, and temporal stability. We calculate the average ranking of each method on each video and rank the results. As a result, each method is assigned a rank of 1–5 on that video.
The final results are shown in Fig. 10. VECNet has 13 results ranked 1, 5 results ranked 2, and a few results ranked 3-5 out of the 20 videos evaluated. When the histograms are compared, it is clear that our methods produce better results for human subjective evaluations across all baselines.
Table 4: Ablation study for investigating the components of the specific modules. For the MFA module, we ablate the alignment with the Fourier transform. For the DIC module, we ablate the single and dual streams.
Model MFA DIC TSR PSNR SSIM
Fourier single dual stage-1 stage-2
(a) 19.03 0.8838
(b) 19.44 0.8905
(c) 19.86 0.8972
(d) 20.11 0.9016
Figure 11: Visual ablation study on the proposed modules.
4.6 Ablation Study and Analysis
We provide a series of ablation studies to evaluate the effectiveness of each component of the proposed method, which are shown in Table 4, Fig. 11. The experiments are conducted on the DIME dataset.
For the MFA module, we remove it and directly fuse the original unaligned multiple frames to learn reflection maps to generate the final result. It can be seen that MFA brings expected quantitative performance improvements. The results of Fig. 11 (b) are more visually appealing with fewer noises and motion blurs than (a), which further demonstrates the effectiveness of MFA. For the DIC unit, we replace the symmetrical exposure stream with a single path setting. It can be seen that the overexposure correction shows little improvement from Fig. 11 (b). For the TSR unit, we split the TSR unit into two separate stages. Comparing Fig. 11 (c) and (d), we can see that our two-stage strategy helps preserve more color and texture information.
In addition, we present the statistical visualization of the results in the feature space. As shown in Fig. 13, after being processed by our method, the underexposure and overexposure representations tend to be intersected together with the corresponding ground truth. It demonstrates the effectiveness of our method for correcting their exposure representations.
Figure 12: The lux curves in the video are enhanced by different methods. The left part is underexposure cases, while the right part is overexposure cases.
Figure 13: t-SNE [25] visualization when taking a testing video (upper part) and testing set (lower part), for example.
@@ 5 Conclusion
We build the first high-quality paired video exposure correction dataset for dynamic real-world scenes with multi-exposures, camera and object motions, and precise spatial alignment. A benchmark dataset is provided for both training and evaluation of video exposure correction methods. Based on the dataset, we propose a method for dealing with the underexposed and overexposed inputs in a dual-stream manner. By utilizing phase alignment and synthesis modules, the exposure of the videos is well corrected and restored. Experiments demonstrate that the proposed method outperforms several image and video methods in low-level tasks adjacent to video exposure correction.